# A Certified Unlearning Approach without Access to Source Data

### v2

@@ 1 Introduction
Machine learning models have achieved remarkable success across a wide range of applications by leveraging large-scale datasets. However, growing concerns about data privacy and regulatory requirements—such as the General Data Protection Regulation (GDPR, 2016), California Consumer Privacy Act (CCPA, 2018), or the Canadian Consumer Privacy Protection Act (CPPA, 2023)—have led to a pressing demand for mechanisms that allow the removal of specific data points from trained models.
Among these mechanisms, certified unlearning has emerged as a cornerstone, formalizing data removal with rigorous guarantees. In contrast to heuristic methods, certified unlearning ensures that the influence of removed data is provably eliminated from the model, offering both privacy compliance and practical utility. This is typically achieved by bounding the statistical discrepancy between a model retrained without the forget data and an approximated model that simulates this retraining. While fully retraining a model from scratch on the retained data is a straightforward and rigorous solution, it becomes computationally prohibitive for large-scale models and frequent deletion requests. Certified unlearning instead provides a practical alternative: it enables efficient, provably correct data removal without exhaustive retraining (Guo et al., 2019; Neel et al., 2021; Sekhari et al., 2021; Chien et al., 2024; Zhang et al., 2024).
Existing certified unlearning methods generally focus on two key aspects. First, they approximate the retrained model using an efficient technique, often leveraging a single-step Newton update (Guo et al., 2019; Sekhari et al., 2021; Zhang et al., 2024). Second, they inject noise based on differential privacy (Dwork, 2006) principles—commonly via a Gaussian mechanism—to ensure that the retrained and unlearned models are statistically indistinguishable. We adopt this single-step update approach in our work as it strikes a balance between efficiency, theoretical grounding, and empirical effectiveness, as supported by prior studies. Despite their promise, these methods often rely on the assumption that the source data remains accessible during unlearning.
@@
This problem is related to zero-shot unlearning, where no information about the true dataset is available during forgetting, beyond the model. While several relevant methods (Foster et al., 2024; Chundawat et al., 2023; Cha et al., 2023; Ahmed et al., 2025) have been proposed, no theoretical guarantees on certified unlearning exist.
@@ 2 Related Works
Certified Unlearning. Certified unlearning has become a key mechanism for data removal, offering rigorous privacy guarantees while avoiding full retraining costs. Existing methods typically rely on single-step Newton updates, influence functions (Guo et al., 2019; Sekhari et al., 2021; Zhang et al., 2024), or projected gradient descent algorithms (Neel et al., 2021; Chien et al., 2024) combined with a randomization mechanism for statistical indistinguishability.
Source-Free Unlearning. Zero-shot machine unlearning (Chundawat et al., 2023) removes data influence using only model weights and the forget set, using noise-based error maximization and gated knowledge transfer. Another method (Cha et al., 2023) employs adversarial sample generation to preserve decision boundaries while applying gradient ascent on forget data. JiT unlearning (Foster et al., 2024) fine-tunes models with perturbed samples to reduce reliance on forget instances. Bonato et al. (Bonato et al., 2025) modify feature vectors to align forgotten data with the nearest incorrect class, using a surrogate dataset in source-free settings. In a concurrent work, Ahmed et al. (Ahmed et al., 2025) approximate the unlearning update by collecting perturbed points around local minima and utilizing only the forget set. This work provides theoretical bounds justifying the methodology, but do not consider formal certified guarantees. Despite some recent work in zero-shot/source-free unlearning, formal certified guarantees remain an open problem. We address this by ensuring such guarantees using a surrogate dataset in the absence of source data.
@@ 3 Preliminaries and Problem Formulation
1. Exact unlearning involves retraining the model from scratch on the retained data $\mathcal{D}_{r}$ (Bourtoule et al., 2021; Ullah et al., 2021; Dukler et al., 2023), which guarantees complete removal of $\mathcal{D}_{u}$’s influence but is often computationally infeasible. This motivated approximate unlearning methods to replicate exact unlearning at a reduced cost.
@@
Most existing certified approximate unlearning methods employ techniques such as influence functions (Guo et al., 2019), second-order Newton updates (Sekhari et al., 2021; Zhang et al., 2024), or other optimization methods (Chien et al., 2024; Neel et al., 2021) to efficiently adjust $\boldsymbol{w}^{\ast}$. By incorporating carefully calibrated randomness, often through a Gaussian mechanism inspired by differential privacy (Dwork, 2006), approximate unlearning ensures that the statistical properties of $\boldsymbol{w}_{r}$ align with those of $\boldsymbol{w}^{\ast}_{r}$, up to a specified certification budget. In these approaches, an upper bound is placed on the norm of the difference between $\boldsymbol{w}^{\ast}_{r}$ and $\boldsymbol{w}_{r}$, which in turn determines the required noise variance. Details of the relation between differential privacy and certified unlearning are provided in Appendix B.
@@
Definition 3.1 ($(\epsilon,\delta)$-Certified Unlearning (Sekhari et al., 2021)).
@@
Definition 3.1 ensures that the updated model $\boldsymbol{w}_{r}$ is statistically indistinguishable from the retrained model $\boldsymbol{w}^{\ast}_{r}$, up to the parameters $(\epsilon,\delta)$. We adopt the unlearning definition from (Sekhari et al., 2021) due to its versatility and practical benefits. Unlike earlier definitions (Ginart et al., 2019), which rely on the unlearning algorithm being inherently randomized even when no deletions occur.
Second-order unlearning. For certified unlearning, $\mathcal{S}(\cdot)$ typically represents detailed information about the true data samples (Guo et al., 2019; Neel et al., 2021; Sekhari et al., 2021; Chien et al., 2024; Zhang et al., 2024). A common approach is to utilize the second-order Newton update, for which $\mathcal{S}(\cdot)$ refers to the Hessian of $\mathcal{D}$, evaluated on the trained model $\boldsymbol{w}^{\ast}$ (Sekhari et al., 2021; Zhang et al., 2024). These approaches follow a general methodology consisting of updating the model with a single-step Newton update,
@@ 4.2 From Theory to Practice
In this section, we first propose an upper bound using Kullback-Leibler (KL) divergence. While total variation distance would be preferable, we use KL for efficiency due to no access to $\mathcal{D}$. Next, we approximate KL without direct access to exact samples by training a model on $\mathcal{D}_{s}$ and utilizing models as conditional probabilities. We sample from input marginal distribution using energy-based modeling to compute the KL. Finally, we estimate the KL between input marginal distributions using the Donsker-Varadhan variational representation (Donsker & Varadhan, 1983). The details of these steps are explained below.
@@
Sampling from input marginal distribution. Inspired by (Grathwohl et al., 2019), we leverage the implicit energy-based model of the trained model $\boldsymbol{w}^{\ast}$ to sample from the approximated input marginal distribution $\hat{\rho}(\boldsymbol{x})$ given by:
@@ 5 Experiments
We systematically evaluate our approach on synthetic and real-world datasets to demonstrate its effectiveness in achieving certified unlearning. Unless otherwise noted, we adopt a linear training model with forget ratio of $0.1$, and an $L_{2}$ regularization constant $\lambda=0.01$. The loss function is assumed to be $\alpha$-strongly convex, $L$-Lipschitz, $\beta$-smooth, and $\gamma$-Hessian Lipschitz. In line with prior works (Koh & Liang, 2017; Wu et al., 2023b, a; Zhang et al., 2024) we tune $\alpha$, $L$, $\beta$, and $\gamma$ for each experimental setting, which preserves theoretical soundness but may lead to approximate certifications. All details about the parameter study and implementation are given under Appendix E.
@@
Performance Metrics. We evaluate the performance using train, test, retain, and forget accuracies on their respective data splits. Additionally, we employ the unlearning-specific membership inference attack (MIA) (Kurmanji et al., 2023), where an accuracy of 50% means the attack can not distinguish whether a specific sample belongs to the forget or test dataset. Relearn time (RT) (Golatkar et al., 2020) measures how many additional training iterations are required to restore the model’s performance on the forgotten data after it is reintroduced. Intuitively, the unlearned model should give high relearn time scores which indicates the model effectively unlearns the forget dataset. Finally, we report the forget score (FS) (Triantafillou et al., 2024), which quantifies how closely the predictions of the unlearned model align with those of a model retrained from scratch. A higher forget score indicates stronger unlearning and higher indistinguishability between retrained and unlearned models.
@@
Real-World Dataset Experiments. We further evaluate our method on CIFAR10 (Krizhevsky et al., 2009), Caltech256 (Griffin et al., 2007), and StanfordDogs (Khosla et al., 2011), by dividing each dataset into an source and a surrogate subset according to a Dirichlet distribution with concentration $\xi$. Lower values of $\xi$ lead to more skewed class splits and thus greater distributional divergences. We show these results in Figure 2(a). We observe that the required noise variance decreases while increasing the concentration parameter $\xi$. We approximate the KL distance between source and surrogate datasets without accessing source data by using Proposition 4.6. We use embeddings from a ResNet18 (He et al., 2016) model, following (Guo et al., 2019).
@@
Additional experiments with different random seeds, corresponding error bars, and evaluations of our heuristic KL approximation—used for noise calibration without accessing the source data—against KL estimates via the Donsker-Varadhan method (with data access) are reported in Appendix F, highlighting the gap between practical estimation and the exact quantity required for certified unlearning.
@@
Mixed-Linear Network Experiments. While the convexity and smoothness assumptions in 4.1 may not hold for general neural networks, there exist practical architectures that satisfy these conditions while retaining strong utility. To this end, we adopt the mixed-linear networks (Golatkar et al., 2021), which linearizes a pre-trained neural network using a first-order Taylor expansion. Specifically, the network output is approximated via its Neural Tangent Kernel (Jacot et al., 2018) formulation, transforming the objective into a convex optimization problem. This approximation allows for efficient and tractable unlearning while preserving much of the model’s predictive performance.
@@
Unlearning Across Model Architectures. To evaluate the generality of our approach, we train three different architectures on CIFAR-10 using a Dirichlet concentration of $\xi$=100: a single linear layer (”L”), a convolutional layer followed by a linear layer (”C+L”), and two convolutional layers with a linear layer (”2C+L”). As shown in Table 6, our method maintains accuracy comparable to others, while keeping the MIA score close to 50%. Also, the RT metric implies that the unlearning succeeded in removing the influence of the forget samples. Finally, for the C+L architecture, the observed FS1 (+), FS1 (-), and FS2 (-) values are 0.08, 0.08, and 0.05, respectively, while for 2C+L, they are lower at 0.04, 0.04, and 0.02. These results reinforce that the introduced noise is essential for the unlearning process.
@@
MNIST-USPS Experiment. To illustrate our contribution in a practical setting, we consider MNIST (Lecun et al., 1998) and USPS (Hull, 1994) datasets and analyze the following cases. First, we train a model on MNIST and apply our unlearning framework by selecting a random forget set from MNIST while using USPS as the surrogate dataset (M $\rightarrow$ U). Second, we reverse the process, training a model on USPS and unlearning with MNIST as the surrogate (U $\rightarrow$ M). As can be seen from Table 7, in both cases the unlearning performance of our method Unlearn (-) is similar to the other methods we are comparing with.
@@ Appendix B Relationship Between Certified Machine Unlearning and Differential Privacy
Differential privacy (Dwork, 2006) and certified unlearning share a common conceptual foundation. Both concepts aim to provide statistical indistinguishability. However, they focus on achieving indistinguishability in fundamentally different aspects of data handling and model behavior.
@@
Certified unlearning often employs DP-inspired techniques like the Gaussian mechanism to achieve guarantees. Specifically, certification requires finding an upper bound for the norm of the difference between a model retrained from scratch and one updated via the unlearning mechanism. This bound enables noise scaling according to the Gaussian mechanism ((Dwork, 2006), App. A) to satisfy $(\epsilon,\delta)$-certified unlearning, ensuring statistical indistinguishability between the unlearned and retrained models (Definition 3.1).
Previous works derive this upper bound based on the unlearning mechanism’s algorithm. Many certified unlearning approaches rely on single-step second-order Newton updates under strong convexity assumptions (Guo et al., 2019; Sekhari et al., 2021; Zhang et al., 2024), where the Hessian is computed over the retain dataset $\mathcal{D}_{r}$ at the model trained on the full dataset $\boldsymbol{w}^{\ast}$.
@@ B.1 Certified Unlearning Using Newton Updates with Exact Data Samples
Certified unlearning methods often utilize statistical information, $\mathcal{S}(\cdot)$, such as the Hessian of the full dataset evaluated at $\boldsymbol{w}^{\ast}$ (Sekhari et al., 2021; Zhang et al., 2024). These methods randomize the unlearning mechanism by adding noise, following the Gaussian mechanism. The upper bound for the norm of the difference between the retrained and unlearned models is derived using retained data samples (Sekhari et al., 2021; Zhang et al., 2024).
Building on this framework, Sekhari et al. (Sekhari et al., 2021) derived an explicit upper bound for the norm difference, leveraging strong convexity and second-order information. This bound provides critical insights into the statistical guarantees of certified unlearning. In the following the upper bound between
Lemma B.2 ((Sekhari et al., 2021) Lemma 3).
@@
The proof can be found in the Supplementary Material of Sekhari et al. (Sekhari et al., 2021) under C.1. ∎
@@
The Algorithm 1 introduces the Gaussian noise to achieve indistinguishability between the model retrained from scratch $\boldsymbol{w}^{\ast}_{r}$ and the model achieved after the unlearning $\widehat{\boldsymbol{w}}_{r}$. Let’s define $(\boldsymbol{w}^{\ast}_{r})^{\prime}=\widehat{\mathcal{U}}(\emptyset,\mathcal{A}(\mathcal{D}_{r}),\mathcal{S}(\mathcal{D}_{r}))=\boldsymbol{w}^{\ast}_{r}+\boldsymbol{n}$ and $\widehat{\boldsymbol{w}}_{r}^{\prime}=\widehat{\mathcal{U}}(\mathcal{D}_{u},\mathcal{A}(\mathcal{D}),\mathcal{S}(\mathcal{D}_{s}))=\widehat{\boldsymbol{w}}_{r}+\boldsymbol{n}$ where the Gaussian noise $\boldsymbol{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ with $\sigma=(\Delta/\epsilon)\sqrt{2\log(1.25/\delta)}$. Following the proof from Dwork et al. ((Dwork, 2006) Theorem A.1), we can prove that for any set $\mathcal{T}\subseteq\mathcal{H}$,

### v3

@@ 1 Introduction
Machine learning models have achieved remarkable success across a wide range of applications by leveraging large-scale datasets. However, growing concerns about data privacy and regulatory requirements—such as the General Data Protection Regulation (gdpr), California Consumer Privacy Act (ccpa), or the Canadian Consumer Privacy Protection Act (cppa)—have led to a pressing demand for mechanisms that allow the removal of specific data points from trained models.
Among these mechanisms, certified unlearning has emerged as a cornerstone, formalizing data removal with rigorous guarantees. In contrast to heuristic methods, certified unlearning ensures that the influence of removed data is provably eliminated from the model, offering both privacy compliance and practical utility. This is typically achieved by bounding the statistical discrepancy between a model retrained without the forget data and an approximated model that simulates this retraining. While fully retraining a model from scratch on the retained data is a straightforward and rigorous solution, it becomes computationally prohibitive for large-scale models and frequent deletion requests. Certified unlearning instead provides a practical alternative: it enables efficient, provably correct data removal without exhaustive retraining (guo_certified_2019; neel_descent--delete_2021; sekhari_remember_2021; chien_langevin_2024; zhang_towards_2024).
Existing certified unlearning methods generally focus on two key aspects. First, they approximate the retrained model using an efficient technique, often leveraging a single-step Newton update (guo_certified_2019; sekhari_remember_2021; zhang_towards_2024). Second, they inject noise based on differential privacy (dwork2006differential) principles—commonly via a Gaussian mechanism—to ensure that the retrained and unlearned models are statistically indistinguishable. We adopt this single-step update approach in our work as it strikes a balance between efficiency, theoretical grounding, and empirical effectiveness, as supported by prior studies. Despite their promise, these methods often rely on the assumption that the source data remains accessible during unlearning.
@@
This problem is related to zero-shot unlearning, where no information about the true dataset is available during forgetting, beyond the model. While several relevant methods (foster_information_2024; chundawat_zero-shot_2023; cha_learning_2023; Ahmed_2025_CVPR) have been proposed, no theoretical guarantees on certified unlearning exist.
@@ 2 Related Works
Certified Unlearning. Certified unlearning has become a key mechanism for data removal, offering rigorous privacy guarantees while avoiding full retraining costs. Existing methods typically rely on single-step Newton updates, influence functions (guo_certified_2019; sekhari_remember_2021; zhang_towards_2024), or projected gradient descent algorithms (neel_descent--delete_2021; chien_langevin_2024) combined with a randomization mechanism for statistical indistinguishability.
Source-Free Unlearning. Zero-shot machine unlearning (chundawat_zero-shot_2023) removes data influence using only model weights and the forget set, using noise-based error maximization and gated knowledge transfer. Another method (cha_learning_2023) employs adversarial sample generation to preserve decision boundaries while applying gradient ascent on forget data. JiT unlearning (foster_information_2024) fine-tunes models with perturbed samples to reduce reliance on forget instances. Bonato et al. (bonato_is_2025) modify feature vectors to align forgotten data with the nearest incorrect class, using a surrogate dataset in source-free settings. In a concurrent work, Ahmed et al. (Ahmed_2025_CVPR) approximate the unlearning update by collecting perturbed points around local minima and utilizing only the forget set. This work provides theoretical bounds justifying the methodology, but do not consider formal certified guarantees. Despite some recent work in zero-shot/source-free unlearning, formal certified guarantees remain an open problem. We address this by ensuring such guarantees using a surrogate dataset in the absence of source data.
@@ 3 Preliminaries and Problem Formulation
1. Exact unlearning involves retraining the model from scratch on the retained data $\mathcal{D}_{r}$ (bourtoule_mu_2021; ullah_machine_2021; dukler2023safe), which guarantees complete removal of $\mathcal{D}_{u}$’s influence but is often computationally infeasible. This motivated approximate unlearning methods to replicate exact unlearning at a reduced cost.
@@
Most existing certified approximate unlearning methods employ techniques such as influence functions (guo_certified_2019), second-order Newton updates (sekhari_remember_2021; zhang_towards_2024), or other optimization methods (chien_langevin_2024; neel_descent--delete_2021) to efficiently adjust $\boldsymbol{w}^{\ast}$. By incorporating carefully calibrated randomness, often through a Gaussian mechanism inspired by differential privacy (dwork2006differential), approximate unlearning ensures that the statistical properties of $\boldsymbol{w}_{r}$ align with those of $\boldsymbol{w}^{\ast}_{r}$, up to a specified certification budget. In these approaches, an upper bound is placed on the norm of the difference between $\boldsymbol{w}^{\ast}_{r}$ and $\boldsymbol{w}_{r}$, which in turn determines the required noise variance. Details of the relation between differential privacy and certified unlearning are provided in Appendix B.
@@
Definition 3.1 ($(\epsilon,\delta)$-Certified Unlearning (sekhari_remember_2021)).
@@
Definition 3.1 ensures that the updated model $\boldsymbol{w}_{r}$ is statistically indistinguishable from the retrained model $\boldsymbol{w}^{\ast}_{r}$, up to the parameters $(\epsilon,\delta)$. We adopt the unlearning definition from (sekhari_remember_2021) due to its versatility and practical benefits. Unlike earlier definitions (ginart2019making), which rely on the unlearning algorithm being inherently randomized even when no deletions occur.
Second-order unlearning. For certified unlearning, $\mathcal{S}(\cdot)$ typically represents detailed information about the true data samples (guo_certified_2019; neel_descent--delete_2021; sekhari_remember_2021; chien_langevin_2024; zhang_towards_2024). A common approach is to utilize the second-order Newton update, for which $\mathcal{S}(\cdot)$ refers to the Hessian of $\mathcal{D}$, evaluated on the trained model $\boldsymbol{w}^{\ast}$ (sekhari_remember_2021; zhang_towards_2024). These approaches follow a general methodology consisting of updating the model with a single-step Newton update,
@@ 4.2 From Theory to Practice
In this section, we first propose an upper bound using Kullback-Leibler (KL) divergence. While total variation distance would be preferable, we use KL divergence for efficiency due to no access to $\mathcal{D}$. Next, we approximate KL divergence without direct access to exact samples by training a model on $\mathcal{D}_{s}$ and utilizing models as conditional probabilities. We sample from input marginal distribution using energy-based modeling to compute the KL divergence. Finally, we estimate the KL divergence between input marginal distributions using the Donsker-Varadhan variational representation (donsker1983asymptotic). The details of these steps are explained below.
@@
Sampling from input marginal distribution. Inspired by (grathwohl_your_2019), we leverage the implicit energy-based model of the trained model $\boldsymbol{w}^{\ast}$ to sample from the approximated input marginal distribution $\hat{\rho}(\boldsymbol{x})$ given by:
@@ 5 Experiments
We systematically evaluate our approach on synthetic and real-world datasets to demonstrate its effectiveness in achieving certified unlearning. Unless otherwise noted, we adopt a linear training model with forget ratio of $0.1$, and an $L_{2}$ regularization constant $\lambda=0.01$. The loss function is assumed to be $\alpha$-strongly convex, $L$-Lipschitz, $\beta$-smooth, and $\gamma$-Hessian Lipschitz. In line with prior works (koh_understanding_2017; wue-edge-23; wu-graph-unlearning-23; zhang_towards_2024) we tune $\alpha$, $L$, $\beta$, and $\gamma$ for each experimental setting, which preserves theoretical soundness but may lead to approximate certifications. All details about the parameter study and implementation are given under Appendix E.
@@
Performance Metrics. We evaluate the performance using train, test, retain, and forget accuracies on their respective data splits. Additionally, we employ the unlearning-specific membership inference attack (MIA) (kurmanji_towards_2023), where an accuracy of 50% means the attack can not distinguish whether a specific sample belongs to the forget or test dataset. Relearn time (RT) (golatkar_eternal_2020) measures how many additional training iterations are required to restore the model’s performance on the forgotten data after it is reintroduced. Intuitively, the unlearned model should give high relearn time scores which indicates the model effectively unlearns the forget dataset. Finally, we report the forget score (FS) (triantafillou2024we), which quantifies how closely the predictions of the unlearned model align with those of a model retrained from scratch. A higher forget score indicates stronger unlearning and higher indistinguishability between retrained and unlearned models.
@@
Real-World Dataset Experiments. We further evaluate our method on CIFAR10 (cifar10), Caltech256 (griffin2007caltech), and StanfordDogs (ssdogs), by dividing each dataset into an source and a surrogate subset according to a Dirichlet distribution with concentration $\xi$. Lower values of $\xi$ lead to more skewed class splits and thus greater distributional divergences. We show these results in Figure 2(a). We observe that the required noise variance decreases while increasing the concentration parameter $\xi$. We approximate the KL divergence between source and surrogate datasets without accessing source data by using Proposition 4.6. We use embeddings from a ResNet18 (resnet18) model, following (guo_certified_2019).
@@
Additional experiments with different random seeds, corresponding error bars, and evaluations of our heuristic KL divergence approximation—used for noise calibration without accessing the source data—against KL divergence estimates via the Donsker-Varadhan method (with data access) are reported in Appendix F, highlighting the gap between practical estimation and the exact quantity required for certified unlearning.
@@
Mixed-Linear Network Experiments. While the convexity and smoothness assumptions in 4.1 may not hold for general neural networks, there exist practical architectures that satisfy these conditions while retaining strong utility. To this end, we adopt the mixed-linear networks (golatkar2021mixed), which linearizes a pre-trained neural network using a first-order Taylor expansion. Specifically, the network output is approximated via its Neural Tangent Kernel (jacot_neural_2018) formulation, transforming the objective into a convex optimization problem. This approximation allows for efficient and tractable unlearning while preserving much of the model’s predictive performance.
@@
Unlearning Across Model Architectures. To evaluate the generality of our approach, we train three different architectures on CIFAR-10 using a Dirichlet concentration of $\xi$=100: a single linear layer (”L”), a convolutional layer followed by a linear layer (”C+L”), and two convolutional layers with a linear layer (”2C+L”). As shown in Table 6, our method maintains accuracy comparable to others, while keeping the MIA score close to 50%. Finally, for the C+L architecture, the observed FS1 (+), FS1 (-), and FS2 (-) values are 0.08, 0.08, and 0.05, respectively, while for 2C+L, they are lower at 0.04, 0.04, and 0.02. These results reinforce that the introduced noise is essential for the unlearning process.
@@
MNIST-USPS Experiment. To illustrate our contribution in a practical setting, we consider MNIST (mnist) and USPS (usps) datasets and analyze the following cases. First, we train a model on MNIST and apply our unlearning framework by selecting a random forget set from MNIST while using USPS as the surrogate dataset (M $\rightarrow$ U). Second, we reverse the process, training a model on USPS and unlearning with MNIST as the surrogate (U $\rightarrow$ M). As can be seen from Table 7, in both cases the unlearning performance of our method Unlearn (-) is similar to the other methods we are comparing with.
@@ Appendix B Relationship Between Certified Machine Unlearning and Differential Privacy
Differential privacy (dwork2006differential) and certified unlearning share a common conceptual foundation. Both concepts aim to provide statistical indistinguishability. However, they focus on achieving indistinguishability in fundamentally different aspects of data handling and model behavior.
@@
Certified unlearning often employs DP-inspired techniques like the Gaussian mechanism to achieve guarantees. Specifically, certification requires finding an upper bound for the norm of the difference between a model retrained from scratch and one updated via the unlearning mechanism. This bound enables noise scaling according to the Gaussian mechanism ((dwork2006differential), App. A) to satisfy $(\epsilon,\delta)$-certified unlearning, ensuring statistical indistinguishability between the unlearned and retrained models (Definition 3.1).
Previous works derive this upper bound based on the unlearning mechanism’s algorithm. Many certified unlearning approaches rely on single-step second-order Newton updates under strong convexity assumptions (guo_certified_2019; sekhari_remember_2021; zhang_towards_2024), where the Hessian is computed over the retain dataset $\mathcal{D}_{r}$ at the model trained on the full dataset $\boldsymbol{w}^{\ast}$.
@@ B.1 Certified Unlearning Using Newton Updates with Exact Data Samples
Certified unlearning methods often utilize statistical information, $\mathcal{S}(\cdot)$, such as the Hessian of the full dataset evaluated at $\boldsymbol{w}^{\ast}$ (sekhari_remember_2021; zhang_towards_2024). These methods randomize the unlearning mechanism by adding noise, following the Gaussian mechanism. The upper bound for the norm of the difference between the retrained and unlearned models is derived using retained data samples (sekhari_remember_2021; zhang_towards_2024).
Building on this framework, Sekhari et al. (sekhari_remember_2021) derived an explicit upper bound for the norm difference, leveraging strong convexity and second-order information. This bound provides critical insights into the statistical guarantees of certified unlearning. In the following the upper bound between
Lemma B.2 ((sekhari_remember_2021) Lemma 3).
@@
The proof can be found in the Supplementary Material of Sekhari et al. (sekhari_remember_2021) under C.1. ∎
@@
The Algorithm 1 introduces the Gaussian noise to achieve indistinguishability between the model retrained from scratch $\boldsymbol{w}^{\ast}_{r}$ and the model achieved after the unlearning $\widehat{\boldsymbol{w}}_{r}$. Let’s define $(\boldsymbol{w}^{\ast}_{r})^{\prime}=\widehat{\mathcal{U}}(\emptyset,\mathcal{A}(\mathcal{D}_{r}),\mathcal{S}(\mathcal{D}_{r}))=\boldsymbol{w}^{\ast}_{r}+\boldsymbol{n}$ and $\widehat{\boldsymbol{w}}_{r}^{\prime}=\widehat{\mathcal{U}}(\mathcal{D}_{u},\mathcal{A}(\mathcal{D}),\mathcal{S}(\mathcal{D}_{s}))=\widehat{\boldsymbol{w}}_{r}+\boldsymbol{n}$ where the Gaussian noise $\boldsymbol{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ with $\sigma=(\Delta/\epsilon)\sqrt{2\log(1.25/\delta)}$. Following the proof from Dwork et al. ((dwork2006differential) Theorem A.1), we can prove that for any set $\mathcal{T}\subseteq\mathcal{H}$,